# Error Correction Code Transformer: From Non-Unified to Unified

### v2

@@ 5.3 Computational Complexity Analysis
Before delving into the computational complexity, let’s first review the unified error correction code Transformer presented in Figure 8, which incorporates $L$ layers of Transformer encoders. The principal computational complexity within each layer is found in the multi-head self-attention module, which is predominantly responsible for matrix multiplication and is defined by the hyperparameters $H$, $d_{k}$ and $d_{l}$. Consequently, the computational complexity of the proposed architecture is given by $O\left(L\times H\times\left(N\times d_{l}\times d_{k}\times H_{d}\right)\right)$, where $N=2n-k$, and $d_{l}\ll N$. Thanks to the sparsity of the parity-check matrix, $H_{d}$ is usually quite small. In our experiment, $H_{d}$ is approximately 0.14, which significantly reduces the computational complexity of the neural network. By contrast, the vanilla Transformer architectures in ECCT and FECCT have complexities of $O\left(L\times H\times\left(N^{2}\times d_{k}\times M_{d}+N\times d_{k}^{2}\right)\right)$ and $O\left(L\times H\times\left(N^{2}\times d_{k}+N\times d_{k}^{2}+N^{2}\times M_{d}\right)\right)$, respectively, where $M_{d}$ represents the mask density in these models.

### v3

@@ 5.3 Computational Complexity Analysis
Before delving into the computational complexity, let’s first review the unified error correction code Transformer presented in Figure 9, which incorporates $L$ layers of Transformer encoders. The principal computational complexity within each layer is found in the multi-head self-attention module, which is predominantly responsible for matrix multiplication and is defined by the hyperparameters $H$, $d_{k}$ and $d_{l}$. Consequently, the computational complexity of the proposed architecture is given by $\mathcal{O}\left(L\times H\times\left(N\times d_{l}\times d_{k}\times H_{d}\right)\right)$, where N=2nkN=2n-k, and $d_{l}\ll N$. Thanks to the sparsity of the parity-check matrix, $H_{d}$ is usually quite small. In our experiment, $H_{d}$ is approximately 0.14, which significantly reduces the computational complexity of the neural network. By contrast, the vanilla Transformer architectures in ECCT and FECCT have complexities of $\mathcal{O}\left(L\times H\times\left(N^{2}\times d_{k}\times M_{d}+N\times d_{k}^{2}\right)\right)$ and $\mathcal{O}\left(L\times H\times\left(N^{2}\times d_{k}+N\times d_{k}^{2}+N^{2}\times M_{d}\right)\right)$, respectively, where $M_{d}$ represents the mask density in these models.